OpenStudy (anonymous):

3w*w=48 What is the value of w? Write the answer in simplest terms.

OpenStudy (anonymous):

3w*w=48 3w^2=48 w^2= 48/3 w^2 = 16 w=square root of 16 that is 4

OpenStudy (anonymous):

thank u im totally lost on this stuff

OpenStudy (anonymous):

you are welcome